### v2

@@ 6 Extensions and Discussions
Building on our earlier work—where we proposed the Sorted-F algorithm and established its constant-factor competitive ratio for this NP-hard problem—we now delve deeper into the problem’s underlying structure. Specifically, we analyze it through the lenses of integer programming (IP) and linear programming (LP), exploring alternative heuristics derived from these frameworks. This investigation not only broadens the theoretical understanding of the problem but also enables a systematic comparison between IP/LP-based approaches and our Sorted-F algorithm, shedding light on their relative strengths and practical trade-offs.
From Jaillet et al. (2025), the optimal schedule can be solved by the following integer programming formulation:
$\text{OPT-IP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (4)
@@
where $\bar{T}$ is the upper bound of the time horizon. The decision variables are $x_{i,t}$. If $x_{i,t}=1$, the request $i$ should start at time $t$. Equation (5) guarantees that each request should only start once. Equation (6) ensures that at any time $t$, the total memory usage in the KV cache does not exceed the memory limit $M$.
Since solving IP is computationally intractable, we consider relaxing the integer variables $x_{i,t}$ to be continuous variables between 0 and $1$, and we have the following relaxed LP formulation:
$\text{Relax-LP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (8)
@@
This relaxation preserves all constraints while making the problem computationally tractable. However, the optimal LP solutions $x_{i,t}^{*}$ typically yield fractional values between 0 and 1, requiring specialized techniques to convert them into actionable scheduling decisions for the discrete original problem. To bridge this gap, we introduce Sorted-LP, an LP-based heuristic that systematically transforms fractional LP solutions into executable schedules.
The algorithm proceeds in three key steps: first, it solves the Relax-LP formulation to obtain fractional assignment variables $x_{i,t}^{*}$. These values are interpreted probabilistically, where each $x_{i,t}^{*}$ represents the likelihood that request $i$ initiates processing at time $t$. Second, for each request $i$, we compute its expected starting time as $y_{i}=\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}$. Finally, Sorted-LP prioritizes requests based on these computed $y_{i}$ values. By sorting all requests in ascending order of $y_{i}$, the algorithm naturally favors requests that the relaxed solution recommends starting earlier. This approach provides a principled way to translate optimization insights from the LP relaxation into a practical scheduling policy. The detailed procedure is formalized in Algorithm 6, which can be found in Appendix 11.1).
While Sorted-LP provides a systematic way to derive schedules from LP solutions, we recognize that further refinement may be possible by incorporating our F-metric framework. To this end, we propose LP-Swap, a hybrid approach that enhances LP-based scheduling with F-metric-guided optimizations. Building on the initial ordering from Sorted-LP, LP-Swap applies local swap operations driven by the quality metric $F(\mathcal{X})$. This hybrid strategy leverages both the global perspective captured by the LP solution and the local optimization capabilities of F-metric refinements. By integrating these complementary approaches, LP-Swap aims to achieve schedule quality beyond what standalone LP solutions can deliver. The complete algorithm pseudocode is provided in Algorithm 7 in Appendix 11.1. We evaluate the effectiveness of the proposed LP-based algorithms using simulated data across a range of distributions in Appendix 11.2, and compare them on the actual dataset in Section 7.
@@ 7 Numerical Experiments
Dataset Overview. To evaluate LLM inference performance under varying input prompt lengths—particularly in scenarios mixing short and long prompts—we combine two publicly available datasets, as no single existing dataset meets this need. The first dataset, introduced by Zheng et al. (2023), contains conversational data from over 210,000 unique IP addresses, collected via the Vicuna demo and Chatbot Arena. This dataset, available at https://huggingface.co/datasets/lmsys/lmsys-chat-1m, consists of short, interactive exchanges with a mean input length of 41.84 tokens (median: 12.00) and a mean output length of 85.35 tokens (median: 43.00), as shown in Figure 5.
Figure 5: Distribution of the number of tokens of input prompt and output response respectively in the data Zheng et al. (2023)
@@
Figure 6: Distribution of the number of tokens of input prompt and output response respectively in the data Cohan et al. (2018)
@@ 8 Supplementary Materials for Section 2
  • Type 2: each request has size $s=1$, output length $o=2$, and there are $Y=M^{1.5}$ such requests.
@@
$\frac{\text{TEL}(\texttt{MC-SF};\mathcal{I})}{\text{TEL}(\Lambda;\mathcal{I})}=\frac{\frac{1}{2}M^{1.5}+3M^{1.25}+M^{2}}{\frac{1}{2}M^{1.5}+3M^{1.25}+6M^{1.5}}\geq\frac{2}{13}M^{0.5}.$
@@
Theorem 8.5
Under the model suggested in Jaillet et al. (2025), suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then minimizing the makespan is NP-hard.
Proof 8.6
Proof of Theorem 8.5 We prove NP-hardness via a reduction from the Partition problem. Consider the following Partition Problem: Given a multiset of integers $X=\{x_{1},\dots,x_{n}\}$ summing to $2T$, does there exist a partition of $X$ into two subsets $S_{1},S_{2}$ such that $\sum_{i\in S_{1}}x_{i}=\sum_{i\in S_{2}}x_{i}=T$?
We construct the following reduction:
  • For each integer $x_{i}$, create a request $i$ with $s_{i}=x_{i}$ and $o_{i}=1$
  • Set memory capacity $M=T$.
We show that the Partition instance has a solution if and only if the constructed instance admits a schedule with makespan 2.
Case 1: Partition Exists. Given subsets $S_{1},S_{2}$ each summing to $T$, the schedule:
  • Batch 1: Process all requests in $S_{1}$ (memory usage = $T$)
  • Batch 2: Process all requests in $S_{2}$ (memory usage = $T$)
completes all requests in makespan 2.
Case 2: No Partition Exists. Any valid schedule must use at least 3 batches because:
  • No single batch can process all requests (total memory $2T>M$)
  • Any two-batch solution would require both batches to have memory exactly $T$, which would constitute a valid partition
[…]
@@ 9 Supplementary Materials for Section 4
$o_{k+1,n_{k+1}}$ $\geq\frac{1}{q_{k+1}}\sum_{j=1}^{q_{k+1}}o_{k+1,j}$
$>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}o_{k,i}$
$>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{2}o_{k,n_{k}}$
$>o_{k,n_{k}}.$ (14)
@@
$o_{k+2,n_{k+2}}$ $\geq\frac{\sum_{l=1}^{q_{k+2}}o_{k+2,l}}{q_{k+2}}$
$>\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{i=1}^{n_{k+1}}o_{k+1,i}}{n_{k+1}}$
$\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{q_{k+1}}$
$\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{\sum_{i=1}^{n_{k}}o_{k,i}}{n_{k}}$
$>4\cdot\frac{1}{2}o_{k,n_{k}}$
@@
Based on Inequalities 14 and 15, for any $i\in\{1,\dots,a_{m}\}$, we can derive the following geometric bound for all batch indices $i\in\{1,\dots,a_{m}\}$:
@@ 10 Supplementary Materials for Section 5
Algorithm 3 Scaled $(1+\epsilon)$-Approximation via Dynamic Programming
[…]
3:Initialize $dp[k][\hat{m}$$\leftarrow\infty$ and $path[k][\hat{m}$$\leftarrow\emptyset$
4:dp[0][0]0dp[0][$\leftarrow 0$
5:for each request $r_{i}\in\mathcal{I}$ do
6:  $\hat{m}_{i}\leftarrow\lfloor(s_{i}+o_{i})/\lambda\rfloor$
7:  for $k=n$ downto 11 do
8:   for $\hat{m}=\hat{M}$ downto $\hat{m}_{i}$ do
9:     if $dp[k-1][\hat{m}-\hat{m}_{i}]+o_{i}<dp[k][\hat{m}$] then
10:      $dp[k][\hat{m}]\leftarrow dp[k-1][\hat{m}-\hat{m}_{i}]+o_{i}$
11:      $path[k][\hat{m}]\leftarrow path[k-1][\hat{m}-\hat{m}_{i}]\cup\{i\}$
12:     end if
13:   end for
14:  end for
15:end for
[…]
17:for $k=1$ to $n$ do
18:  for $\hat{m}=0$ to $\hat{M}$ do
19:   $F\leftarrow dp[k][\hat{m}]/k^{2}$
20:   if $F<F^{*}$ then
21:     $F^{*}\leftarrow F$
22:     $\mathcal{X}^{*}\leftarrow path[k][\hat{m}$]
23:   end if
24:  end for
25:end for
26:return $\mathcal{X}^{*}$
Algorithm 4 Local Swap Search Heuristic
@@ 11.1 Pseudocodes for LP-Based Methods
1:Solve Relax-LP and obtain optimal solutions $x_{i,t}^{*}$ for all $i\in[n],t\in[\bar{T}]$
@@
5:Sort all requests $r_{i}\in\mathcal{I}$ in ascending order of $y_{i}$ to obtain $\mathcal{I}^{\prime}$
@@
7:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
8:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
9:  for $r_{i}\in\mathcal{R}^{(t)}$ in order of $\mathcal{I}^{\prime}$ do
10:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
11:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
12:   else
13:     Break
14:   end if
15:  end for
@@
1:Solve Relax-LP and obtain optimal solutions $x_{i,t}^{*}$ for all $i\in[n],t\in[\bar{T}]$
2:for each $i\in[n]$ do
3:  Compute $y_{i}\leftarrow\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}$
4:end for
5:[$\mathcal{I}^{\prime}\leftarrow$[]
6:while $\mathcal{I}$ do
7:  Sort all requests $r_{i}\in\mathcal{I}$ in ascending order of $y_{i}$
8:  $\mathcal{X}_{\text{init}}\leftarrow\emptyset,\ mem\leftarrow 0$
9:  for each $r\in\mathcal{I}$ in sorted order do
10:   if $mem+s_{r}+o_{r}\leq M$ then
11:     $\mathcal{X}_{\text{init}}\leftarrow\mathcal{X}_{\text{init}}\cup\{r\}$
12:     $mem\leftarrow mem+s_{r}+o_{r}$
13:   end if
14:  end for
15:  $\mathcal{X}^{*}\leftarrow\text{LocalSwap}(\mathcal{X}_{\text{init}},\mathcal{I},M)$
16:  Sort $r_{i}\in\mathcal{X}^{*}$ in ascending order of $o_{i}$
17:  $\mathcal{I}^{\prime}\leftarrow\mathcal{I}^{\prime}+\mathcal{X}$, $\mathcal{I}\leftarrow\mathcal{I}-\mathcal{X}$
18:end while
@@
20:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
21:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
22:  for $r_{i}\in\mathcal{R}^{(t)}$ in sorted order do
23:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
24:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
25:   else
26:     Break
27:   end if
28:  end for

### v3

@@ 6 Extensions and Discussions
Building on our earlier work—where we proposed the Sorted-F algorithm and established its constant-factor competitive ratio for this NP-hard problem—we now discuss alternative formulations and heuristics based on integer programming (IP) and linear programming (LP). The goal of this section is twofold: (i) to use IP/LP to further illuminate the structure of the problem (e.g., via relaxations and lower bounds), and (ii) to derive practical LP-guided scheduling heuristics that serve as complementary baselines in our numerical evaluation. We emphasize that the LP-based methods below, Sorted-LP and LP-Swap, are heuristics: we do not provide worst-case approximation/competitive-ratio guarantees for their rounded schedules, and the integrality gap of the relaxation is not characterized here.
From Jaillet et al. (2025), the offline optimum can be formulated as the following integer program:
$\text{Optimal-IP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (7)
@@
where $\bar{T}$ is an upper bound on the time horizon. The decision variables are $x_{i,t}$. If $x_{i,t}=1$, request $i$ starts at time $t$. Constraint (8) enforces that each request starts exactly once, and constraint (9) ensures that at any time $t$, the total KV-cache memory usage does not exceed $M$.
Since solving Optimal-IP is computationally intractable, we relax the integrality constraints and obtain the following LP:
$\text{Relaxed-LP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (11)
@@
The Relaxed-LP objective value provides a valid lower bound on Optimal-IP and thus on the offline optimal latency. However, optimal LP solutions $x_{i,t}^{*}$ are typically fractional, and converting them into a feasible discrete schedule requires rounding or other heuristics. In what follows, we introduce two LP-guided scheduling heuristics, Sorted-LP and LP-Swap, whose effectiveness is assessed empirically (Appendix 11.2 and Section 7). We do not claim provable guarantees for these rounded schedules; in particular, the integrality gap of Relaxed-LP and the approximation quality of the rounding procedures are left for future work.
Sorted-LP: LP-induced ordering heuristic.
The algorithm proceeds in three steps. First, it solves Relaxed-LP to obtain an optimal fractional solution $x_{i,t}^{*}$. Second, for each request $i$, it computes the LP-induced average expected start time
$y_{i}\;=\;\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}.$
Finally, Sorted-LP sorts requests by increasing $y_{i}$ and then executes them using the same feasible admission rule as in Phase 2 of Sorted-F. Intuitively, smaller $y_{i}$ indicates that the LP relaxation “prefers” starting request $i$ earlier; we use this as an ordering signal rather than a guarantee. The detailed procedure is formalized in Algorithm 6 in Appendix 11.1.
LP-Swap: LP ordering + F-metric refinement.
While Sorted-LP provides a principled ordering signal from the LP relaxation, it may be further improved by incorporating our F-metric framework. To this end, we propose LP-Swap, a hybrid approach that starts from the Sorted-LP-induced ordering and then applies F-metric-guided local swap refinements when forming batches. This hybrid strategy leverages both the global perspective captured by the LP relaxation and the local optimization capability of the swap procedure. The complete pseudocode is provided in Algorithm 7 in Appendix 11.1. We evaluate the effectiveness of these LP-based heuristics using simulated data in Appendix 11.2 and compare them on the real dataset in Section 7.
@@ 7 Numerical Experiments
Dataset Overview. To evaluate LLM inference performance under varying input prompt lengths—particularly in scenarios mixing short and long prompts—we combine two publicly available datasets, as no single existing dataset meets this need. The first dataset, introduced by Zheng et al. (2023a), contains conversational data from over 210,000 unique IP addresses, collected via the Vicuna demo and Chatbot Arena. This dataset, available at https://huggingface.co/datasets/lmsys/lmsys-chat-1m, consists of short, interactive exchanges with a mean input length of 41.84 tokens (median: 12.00) and a mean output length of 85.35 tokens (median: 43.00), as shown in Figure 6.
Figure 6: Upper: Distribution of the number of tokens of input prompt and output response respectively in the data Zheng et al. (2023a). Lower: Distribution of the number of tokens of input prompt and output response respectively in the data Cohan et al. (2018).
@@ 8 Supplementary Materials for Section 2
  • Type 2: each request has size $s=1$, output length $o=2$, and there are $Y=M^{1.25}$ such requests.
@@
$\frac{\text{TEL}(\texttt{MC-SF};\mathcal{I})}{\text{TEL}(\Lambda;\mathcal{I})}=\frac{M^{1.75}+\frac{7}{2}M^{1.5}+M^{1.25}+\frac{1}{2}M}{\frac{7}{2}M^{1.5}+7M^{1.25}+\frac{1}{2}M}\geq\frac{1}{11}M^{0.25}.$
@@ 9 Supplementary Materials for Section 4
$o_{k+1,n_{k+1}}$ $\geq\frac{1}{q_{k+1}}\sum_{j=1}^{q_{k+1}}o_{k+1,j}>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}o_{k,i}>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{2}o_{k,n_{k}}>o_{k,n_{k}}.$ (17)
@@
$o_{k+2,n_{k+2}}$ $\geq\frac{\sum_{l=1}^{q_{k+2}}o_{k+2,l}}{q_{k+2}}>\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{i=1}^{n_{k+1}}o_{k+1,i}}{n_{k+1}}\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{q_{k+1}}$
$\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{\sum_{i=1}^{n_{k}}o_{k,i}}{n_{k}}>4\cdot\frac{1}{2}o_{k,n_{k}}$
@@
Based on Inequalities 17 and 18, for any $i\in\{1,\dots,a_{m}\}$, we can derive the following geometric bound for all batch indices $i\in\{1,\dots,a_{m}\}$:
@@
[…]
Proof 9.2
Proof of Lemma 4.11
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\mathrm{o}\left(\sum_{k=b_{m}+1}^{\infty}n_{k}\right)\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\sum_{k=b_{m-1}+1}^{b_{m}}n_{k}\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
[…]
Proof 9.3
Proof of Lemma 4.18:
$\mathrm{TEL}(\textup{{Optimal}}_{\mathrm{align}};\mathcal{I})$ $=\sum_{h=1}^{\infty}\left(\bar{o}_{h}\cdot\sum_{j=h+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=\sum_{g=1}^{\infty}\sum_{h=d_{g-1}+1}^{d_{g}}\left(\bar{o}_{h}\cdot\left(\mathrm{o}\left(\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{j=d_{g}+1}^{\infty}n_{j}\right)\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$\leq 6\cdot\left(\sum_{g=1}^{\infty}\left(\delta_{g}\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}+\sum_{r_{\ell}\in\mathcal{Z}_{g}}(p_{\ell}-t_{g})\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}\right)$
$=6\cdot\mathrm{TEL}(\textup{{Optimal}};\mathcal{I}).$
@@ 10 Supplementary Materials for Section 5
Algorithm 3 Local Swap Search
@@
[…]
1:Request set $\mathcal{I}$ with prefill sizes $\{s_{i}\}$ and predicted intervals $\{[\ell_{i},u_{i}]\}$; memory capacity $M$
2:Schedule minimizing total end-to-end latency
3:Initialization: Set $\tilde{o}_{i}\leftarrow\ell_{i}$ for all $i\in\mathcal{I}$ $\triangleright$ Optimistic initialization
4:
5:Phase 1: Batch Construction with Lower Bounds
6:Run Phase 1 of Sorted-F (Algorithm 1) using $\tilde{o}_{i}$ in place of $o_{i}$
7:Let $\mathcal{I}^{\prime}$ be the resulting ordered request sequence
8:
9:Phase 2: Adaptive Scheduling Execution
[…]
11:while $\mathcal{R}^{(t)}\neq\emptyset$ or $\mathcal{V}^{(t)}\neq\emptyset$ do
12:  $t\leftarrow t+1$
13:  for $r_{i}\in\mathcal{V}^{(t-1)}$ do $\triangleright$ Update running estimates
14:   $\tilde{o}_{i}\leftarrow\max\{\tilde{o}_{i},\,t-p_{i}\}$
15:  end for
16:  Remove completed requests: $\mathcal{V}^{(t)}\leftarrow\{r_{i}\in\mathcal{V}^{(t-1)}:t-p_{i}<\tilde{o}_{i}\text{ or }r_{i}\text{ not yet terminated}\}$
17:
18:  Admission Step: Attempt to add pending requests to active set
19:  Run admission logic of Phase 2 of Sorted-F using current $\{\tilde{o}_{i}\}$ values
20:  Let $\mathcal{U}^{(t)}$ be the set of newly admitted requests; set $p_{i}\leftarrow t$ for $r_{i}\in\mathcal{U}^{(t)}$
21:
22:  Eviction Step: Resolve memory overflow if necessary
23:  while $M(\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)},\,\{\tilde{o}_{j}\})>M$ do $\triangleright$ Ordered eviction from $\mathcal{A}_{\min}$
24:   $r^{*}\leftarrow\arg\min_{r_{i}\in\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}}\tilde{o}_{i}$ $\triangleright$ Evict request closest to completion
25:   $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t)}\setminus\{r^{*}\}$;   $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}\setminus\{r^{*}\}$
26:   Return $r^{*}$ to pending queue $\mathcal{R}^{(t)}$ with updated $\tilde{o}_{r^{*}}$
27:  end while
28:
29:  $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\setminus\mathcal{U}^{(t)}$
30:  $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}$
[…]
32:end while
@@ 11.1 Pseudocodes for LP-Based Methods
1:Solve Relaxed-LP and obtain an optimal fractional solution $x_{i,t}^{*}$ for all $i\in[n],t\in[\bar{T}]$
@@
5:Sort all requests $r_{i}\in\mathcal{I}$ in ascending order of $y_{i}$ to obtain an ordered list $\mathcal{I}^{\prime}$
@@
7:while $\mathcal{R}^{(t)}\neq\emptyset\ \textbf{or}\ \mathcal{V}^{(t)}\neq\emptyset$ do
8:  $t\leftarrow t+1$
9:  $\mathcal{U}^{(t)}\leftarrow\emptyset$
10:  $\mathcal{V}^{(t)}\leftarrow\{r_{j}\in\mathcal{V}^{(t-1)}:p_{j}+o_{j}\geq t\}$
11:  for $r_{i}\in\mathcal{R}^{(t-1)}$ in order do
12:   tentatively set $p_{i}\leftarrow t$
13:   $\widehat{\mathcal{U}}\leftarrow\mathcal{U}^{(t)}\cup\{r_{i}\}$
14:   if $M(\mathcal{V}^{(t)}\cup\widehat{\mathcal{U}},\,p_{j}+o_{j})\leq M,\ \forall r_{j}\in\mathcal{V}^{(t)}\cup\widehat{\mathcal{U}}$ then
15:     $\mathcal{U}^{(t)}\leftarrow\widehat{\mathcal{U}}$
16:   else
17:     break
18:   end if
19:  end for
20:  $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}\setminus\mathcal{U}^{(t)}$
21:  $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}$
@@
1:Solve Relaxed-LP and obtain $x^{*}_{i,t}$ for all $i\in[n],\,t\in[\bar{T}]$; set $y_{i}\leftarrow\sum_{t\in[\bar{T}]}t\,x^{*}_{i,t}$ for all $i\in[n]$
2:Helper: $\textsc{Feasible}(\mathcal{S})\equiv\big(M(\mathcal{S},\,p_{j}+o_{j})\leq M,\ \forall r_{j}\in\mathcal{S}\big)$
3:[$\mathcal{I}^{\prime}\leftarrow$[]
4:while $\mathcal{I}\neq\emptyset$ do
5:  Sort $\mathcal{I}$ by increasing $y_{i}$
6:  $\mathcal{X}_{\mathrm{init}}\leftarrow\emptyset,\ mem\leftarrow 0$
7:  for each $r\in\mathcal{I}$ in this order do
8:   if $mem+s_{r}+o_{r}\leq M$ then
9:     $\mathcal{X}_{\mathrm{init}}\leftarrow\mathcal{X}_{\mathrm{init}}\cup\{r\}$; $mem\leftarrow mem+s_{r}+o_{r}$
10:   end if
11:  end for
12:  $\mathcal{X}^{*}\leftarrow\textsc{LocalSwap}(\mathcal{X}_{\mathrm{init}},\mathcal{I},M)$
13:  Sort $\mathcal{X}^{*}$ by nondecreasing $o_{i}$
14:  Append $\mathcal{X}^{*}$ to $\mathcal{I}^{\prime}$
15:  $\mathcal{I}\leftarrow\mathcal{I}\setminus\mathcal{X}^{*}$
16:end while
@@
18:while $\mathcal{R}^{(t)}\neq\emptyset\ \textbf{or}\ \mathcal{V}^{(t)}\neq\emptyset$ do
19:  $t\leftarrow t+1$; $\mathcal{U}^{(t)}\leftarrow\emptyset$; $\mathcal{V}^{(t)}\leftarrow\{r_{j}\in\mathcal{V}^{(t-1)}:p_{j}+o_{j}\geq t\}$
20:  for $r_{i}\in\mathcal{R}^{(t-1)}$ in order do
21:   tentatively set $p_{i}\leftarrow t$
22:   if $\textsc{Feasible}(\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}\cup\{r_{i}\})$ then
23:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}\cup\{r_{i}\}$
24:   else
25:     break
26:   end if
27:  end for
28:  $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}\setminus\mathcal{U}^{(t)}$; $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t)}\cup\mathcal{U}^{(t)}$